OpenStudy (anonymous):

A pharmaceutical company sells bottles of 500 calcium tablets in two dosages: 250 milligram and 500 milligram. Last month, the company sold 2,200 bottles of 250-milligram tablets and 1,800 bottles of 500-milligram tablets. The total sales revenue was $39,200. The sales team has targeted sales of $44,000 for this month, to be achieved by selling of 2,200 bottles of each dosage. Assuming that the prices of the 250-milligram and 500-milligram bottles remain the same, what's the price of a 250-milligram bottle ? and what's the price of a 500-milligram bottle ?

OpenStudy (texaschic101):

t = 250 mg f = 500 mg 2200t + 1800f = 39,200 --->(multiply by -1) 2200t + 2200f = 44,000 ---------------------- -2200t - 1800f = - 39,200 (result of multiplying by -1) 2200t + 2200f = 44000 ----------------------add 400f = 4800 f = 4800/400 f = 12 2200t + 1800f = 39,200 2200t + 1800(12) = 39,200 2200t + 21600 = 39200 2200t = 39200 - 21600 2200t = 17600 t = 17600/2200 t = 8 check.. 2200t + 2200f = 44000 2200(8) + 2200(12) = 44000 17600 + 26400 = 44000 44000 = 44000 (correct) 250 mg bottles (t) cost $8 per bottle 500 mg bottles (f) cost $12 per bottle

OpenStudy (anonymous):

Thanksss